OpenStudy (anonymous):

Which of the following shows the square completed in the this equation? x^2+7x+2=0 A. (x+7x)^2+2=0 B. (x+3.5)^2=10.25 C. (x+2.6)^2=4.76

OpenStudy (across):

x^2+7x+2=0 x^2+7x=-2 x^2+7x+12.25=10.25 (x+3.5)^2=10.25